OpenStudy (anonymous):

Solve for m. 2 – 3m = -16

OpenStudy (gabebae):

I'll help

OpenStudy (anonymous):

3m = -18 m = -18/3 m = -6

OpenStudy (gabebae):

2-3m=-16 -2 -2 ---------- -3m = -18 - - -3 -3 ------------ m = 6

OpenStudy (anonymous):

Metal me. plz.

OpenStudy (anonymous):

It was 6.

OpenStudy (gabebae):

You're welcome

OpenStudy (anonymous):

I didn't say thanks yet! Thanks.

OpenStudy (anonymous):

looks like I had a sign error this time hehe.